# ReCap: Better Gaussian Relighting with Cross-Environment Captures

### v1

@@ Abstract
Accurate 3D objects relighting in diverse unseen environments is crucial for realistic virtual object placement. Due to the albedo-lighting ambiguity, existing methods often fall short in producing faithful relights. Without proper constraints, observed training views can be explained by numerous combinations of lighting and material attributes, lacking physical correspondence with the actual environment maps used for relighting. In this work, we present ReCap, treating cross-environment captures as multi-task target to provide the missing supervision that cuts through the entanglement. Specifically, ReCap jointly optimizes multiple lighting representations that share a common set of material attributes. This naturally harmonizes a coherent set of lighting representations around the mutual material attributes, exploiting commonalities and differences across varied object appearances. Such coherence enables physically sound lighting reconstruction and robust material estimation — both essential for accurate relighting. Together with a streamlined shading function and effective post-processing, ReCap outperforms the leading competitor by $3.4$ dB in PSNR on an expanded relighting benchmark.
Figure 1: ReCap shows a significant advantage in reconstructing environment maps with accurate tones and color fidelity. Both qualitative and quantitative assessments show that ReCap achieves more realistic and consistent relighting results under a range of unseen lighting conditions. See the project page.
@@ 1 Introduction
Inspired by photometric appearance modeling [32, 24, 43] which estimates surface properties from object appearances under varied lighting, we introduce additional photometric supervision to address the albedo-lighting ambiguity. While traditional approaches rely on controlled lighting setups like light stages [11] and collocated lights [42] to provide supervision through known light directions and/or intensities, such controlled capture is impractical for real-world applications. Instead, we propose ReCap to leverage object captures across unknown lighting conditions, modeling light-dependent appearances with multiple environment maps that share a common Gaussian model. Conceptually, this resembles multi-task learning, where the learned environment maps act as task heads querying a shared material representation for varied object appearances. In this case, the “querying” is done by the physically-based shading function, which facilitates accurate separation of material and lighting. This joint optimization fosters robust material estimation across diverse environments while maintaining internal consistency and coherence, simultaneously supporting precise lighting reconstruction.
@@
Current relighting evaluations are limited in scope, focusing primarily on diffuse surfaces [19, 49]. For a more comprehensive assessment, we re-rendered $13$ objects from NeRF [27] and RefNeRF [38] featuring both diffuse and specular surfaces. As shown in Fig. 1, experiments on the expanded benchmark confirm the effectiveness of ReCap in producing robust and more realistic relighting results. Codes and datasets will be made publicly available here.
@@
Figure 2: The proposed ReCap training framework. Compared to original 3DGS [22], each Gaussian is augmented with 3 extra material attributes. Given $k$ sets of object appearances from unknown lighting conditions as input, $k$ learnable environment maps are instantiated. Gaussian color is pre-computed according to the shading function in the world space based on environment queries and material properties. 2D images are rasterized with standard Gaussian splatting and used for loss computation. $\mathcal{L}_{\text{image}}$: image reconstruction loss from [22]. Additional loss terms for material and geometry are not shown.
@@ 2 Related Work
Novel View Synthesis (NVS).
Pure NVS techniques focus on reproducing the scene appearance under its original environment. NeRF [27] and its follow-ups [2, 3, 4] achieve remarkable NVS quality with implicit radiance fields and volume rendering. Although great progress has been made in NeRF acceleration [14, 28, 41] leveraging voxel grids or hash tables, their rendering speeds ($\sim$10 fps) are still far from interactive. Building on progress in differentiable point-based rendering [45, 23], 3DGS [22] recently introduced an explicit Gaussian representation with an efficient tile-based rasterizer, delivering photo-realistic rendering at impressive frame rates ($\sim$100 fps). This makes 3DGS well-suited for high-quality real-time relighting.
Relighting.
The relighting task traditionally involves editing lighting within fixed views [12, 35, 42]. The term has since broadened to encompass novel view relighting [43, 20], which aims to generate images from both new viewpoints and under new lighting conditions. With the advent of differentiable neural rendering, reconstruction-based methods [7, 46, 48, 37] have found a promising direction to combine physically-based rendering (PBR) with neural representations for editable and realistic rendering. Closely related to our work, PBR is also incorporated by 3D Gaussians, which enable Gaussian relighting. 3DGS-DR [44] and GShader [18] both specialize in specular objects and propose customized shading functions. R3DG [15] further incorporates ray tracing for indirect illumination, while GS-IR [25] relies on baked occlusion maps. With only single environment captures as input, these Gaussian relighting methods face challenges in correctly decoupling lighting from material properties.
Albedo-lighting Disambiguation.
Recovering reliable illumination and albedo from object appearances is an ill-posed problem [34] in inverse rendering. It is often simplified under the assumption of controllable lighting [6, 30], known geometry [31], or when regularized with strong priors [36, 13] in domain-specific tasks such as relighting of human faces. For more general cases, object captures from varied lighting, known as photometric images, can provide important visual cues. Several NeRF-based relighting methods have taken advantage of such captures. NeRV [37] requires multiple known lighting environments, while TensoIR [19] internalizes lighting as an additional dimension in their factorized tensorial representation. NeRD [7] models lighting as spherical Gaussians and optimizes separate lighting for every input image. In contrast, we use object captures from unknown lighting conditions, with learnable lighting representations optimized independently for each scene as environment maps.
@@ 3.2 Disambiguate Split Sum Approximation
$\begin{split}L_{out}(\mathbf{x},\mathbf{v})&=\underbrace{E_{d}(\mathbf{n})(1-m)\mathbf{b}}_{diffuse}\\ &+\underbrace{E_{s}(\mathbf{n},\mathbf{v})\left[F_{0}\beta_{1}(r,\mathbf{n},\mathbf{v})+\beta_{2}(r,\mathbf{n},\mathbf{v})\right]}_{specular}\\ \end{split}$ (6)
@@ 3.3 Lighting Representation
While increasing the number of environments can enhance the decoupling as later shown in Sec. 4.3, it also requires additional effort to capture real objects across multiple scenes. As a proof of concept, we limit our investigation to a dual-environment setup for the remainder of the text.
@@
3.4 Physically Appropriate Post-processing
Standard HDR maps typically represent radiance in linear RGB space when used as lighting sources in rendering [10]. To properly apply them for relighting, the learned lighting values need a clear physical interpretation. Current methods make diverse design choices during training, which jointly determine the interpretation of the learned environment maps and how new HDR maps should be processed. For example, if the range of the learned values is limited between $[0,1]$ and gamma correction is applied after the shading equation, the learned environment map is interpreted as being in a linear low dynamic range (LDR) space. In Tab. 1, we analyze three major factors that impact relighting performance. Gamma correction is found to be essential for relighting, which is expected, since the shading function models linear light transport. However, it is often overlooked [15, 44], as its omission still provides good NVS results. Complex tone-mappers are common in graphics and are widely adopted [37], but they perform worse than simple clipping, likely due to the introduced non-linearity that hinders optimization.
Striking a balance between relighting and NVS, we constrain our environment maps to be positive, applying simple clipping followed by gamma correction. This ensures a linear HDR lighting space, where new environment maps can be applied without modification. Together with cross-environment decoupling, our learned environment maps are closer to the actual HDR maps in distribution. This removes the need for normalizing relit images [26] to match the average colors of ground-truth images (which are not available in practice) when comparing relighting results.
@@ 3.5 Geometry Estimation
Table 2: Relighting results across different environment maps. ReCap trains on 100 views each from “courtyard” and “bridge”, while other methods use 200 views from “bridge” only. The * symbol indicates learned environment, equivalent to NVS. Note that 3DGS-DR’s good performance on “bridge” reflects its tendency to regurgitate training views. Best results are highlighted in bold.
Unseen Envs Training Envs NVS
night snow sunset interior fireplace forest average courtyard bridge bridge*
@@
GShader [18] 22.33 18.14 19.26 21.87 21.03 23.04 20.95 20.68 26.04 32.92
3DGS-DR [44] 20.88 18.92 23.53 21.00 19.52 25.19 21.51 20.58 33.19 35.50
GS-IR [25] 24.87 21.95 21.15 20.18 24.27 20.47 22.15 20.02 20.72 29.95
R3DG [15] 23.20 22.05 22.93 19.65 22.62 22.26 22.12 20.22 22.25 30.44
ReCap (ours) 26.68 24.56 26.56 24.34 25.51 25.45 25.52 25.51 26.97 32.03
@@
GShader [18] 0.899 0.889 0.900 0.907 0.869 0.907 0.895 0.893 0.950 0.969
3DGS-DR [44] 0.892 0.897 0.931 0.909 0.858 0.910 0.900 0.894 0.971 0.978
GS-IR [25] 0.900 0.896 0.892 0.878 0.876 0.863 0.884 0.871 0.879 0.938
R3DG [15] 0.889 0.914 0.920 0.865 0.867 0.879 0.889 0.871 0.879 0.959
ReCap (ours) 0.932 0.938 0.951 0.929 0.903 0.927 0.930 0.942 0.951 0.965
@@
GShader [18] 0.086 0.115 0.095 0.091 0.106 0.089 0.102 0.097 0.061 0.043
3DGS-DR [44] 0.077 0.089 0.070 0.086 0.101 0.081 0.084 0.095 0.040 0.033
GS-IR [25] 0.104 0.102 0.105 0.112 0.112 0.114 0.108 0.122 0.113 0.071
R3DG [15] 0.104 0.088 0.081 0.126 0.113 0.103 0.108 0.113 0.113 0.053
ReCap (ours) 0.062 0.070 0.057 0.076 0.076 0.071 0.069 0.065 0.058 0.047
@@ 4.1 Implementation Details
We construct a more comprehensive RelightObj dataset by relighting 8 general objects from NeRF Synthetic Dataset [27] and 5 shiny objects from the Shiny Blender Dataset [38] under 8 different HDR maps. Each scene includes 200 training views and 200 test views, with identical camera poses across scenes. For a consistent relighting comparison, all methods are trained with captures from a single environment, except ReCap, which additionally uses a second environment. To ensure fairness in the number of training views, all methods use 200 views from the primary training environment, while ReCap uses 100 views from each of the two training environments. Relighting performance will be judged on unseen environments.
Baseline and Metrics.
@@
We compare our method against the following baselines: a) GShader [18]: A Gasussian rendering method utilizing a customized shading function with a residual color term. b) GS-IR [25]: An inverse rendering approach using view-space shading with a microfacet BRDF model and a baked occlusion map. c) 3DGS-DR [44]: A Gasussian rendering method leveraging a customized shading function and normal propagation in view-space. d) R3DG [15]: A Gaussian relighting technique that incorporates ray tracing and a microfacet BRDF model. Following these works, rendering and relighting quality are evaluated using PSNR, SSIM [40], and LPIPS [47].
Training and Testing.
@@
All experiments are conducted on an Nvidia RTX 3090 graphics card. We optimize our models using the Adam optimizer over 30,000 iterations. All methods are retrained on our dataset using their official repositories with recommended settings, results are reported from the retrained models. Since GShader [18] and 3DGS-DR [44] do not provide official relighting code, we pre-processed the HDR maps according to their design choices, with details in the supplementary material.
@@
Figure 6: Qualitative comparison of relighting results across different environment maps. Specular objects provide more information for lighting reconstruction, leading to better albedo-lighting decoupling hence more truthful relighting results. Refer to the supplementary for more visuals.
Table 3: Object-wise relighting performance. The average is reported over unseen environment maps. ReCap shows the best overall performance for both diffuse and specular objects, while performing comparably on objects with shadows or translucent materials. Best results are highlighted in bold.
Relit NeRF Synthetic Relit Shiny Blender
chair drums ficus hotdog lego ship material mic helmet coffee musclecar toaster teapot
[…]
GShader [18] $20.83$ $19.69$ $25.56$ $18.80$ $18.85$ $17.10$ $20.03$ $27.86$ $19.07$ $\mathbf{20.28}$ $21.51$ $18.55$ $24.14$
3DGS-DR [44] $21.25$ $21.82$ $25.78$ $19.34$ $24.37$ $18.98$ $20.01$ $28.02$ $20.33$ $18.56$ $19.47$ $17.35$ $24.34$
GS-IR [25] $21.87$ $22.37$ $26.75$ $21.74$ $19.89$ $\mathbf{21.18}$ $22.26$ $27.25$ $19.31$ $18.60$ $23.39$ 16.79 26.52
R3DG [15] 23.67 21.23 25.60 24.38 22.21 20.50 21.55 27.67 21.17 16.27 22.74 16.66 23.89
ReCap (ours) 26.63 24.77 28.24 23.25 24.44 20.12 25.83 31.44 23.55 19.78 29.14 25.05 29.49
[…]
GShader [18] 0.914 0.891 0.951 0.885 0.866 0.774 0.871 0.968 0.889 0.931 0.898 0.824 0.973
3DGS-DR [44] 0.920 0.911 0.958 0.882 0.926 0.771 0.875 0.970 0.885 0.921 0.881 0.824 0.971
GS-IR [25] 0.905 0.901 0.951 0.886 0.856 0.783 0.873 0.958 0.850 0.896 0.896 0.774 0.967
R3DG [15] 0.924 0.895 0.952 0.904 0.888 0.775 0.864 0.962 0.865 0.877 0.900 0.784 0.968
ReCap (ours) 0.951 0.937 0.972 0.912 0.918 0.817 0.921 0.978 0.920 0.928 0.944 0.910 0.983
[…]
GShader [18] 0.078 0.068 0.038 0.133 0.099 0.187 0.089 0.024 0.134 0.142 0.073 0.157 0.038
3DGS-DR [44] 0.055 0.056 0.029 0.110 0.055 0.173 0.081 0.020 0.116 0.144 0.075 0.143 0.033
GS-IR [25] 0.068 0.069 0.040 0.133 0.096 0.189 0.093 0.032 0.178 0.164 0.083 0.220 0.041
R3DG [15] 0.060 0.068 0.037 0.105 0.084 0.188 0.103 0.033 0.171 0.157 0.078 0.210 0.038
ReCap (ours) 0.042 0.042 0.023 0.091 0.065 0.167 0.063 0.016 0.094 0.128 0.047 0.088 0.027

### v2

@@ Abstract
11footnotetext: Corresponding author
Accurate 3D objects relighting in diverse unseen environments is crucial for realistic virtual object placement. Due to the albedo-lighting ambiguity, existing methods often fall short in producing faithful relights. Without proper constraints, observed training views can be explained by numerous combinations of lighting and material attributes, lacking physical correspondence with the actual environment maps used for relighting. In this work, we present ReCap, treating cross-environment captures as multi-task target to provide the missing supervision that cuts through the entanglement. Specifically, ReCap jointly optimizes multiple lighting representations that share a common set of material attributes. This naturally harmonizes a coherent set of lighting representations around the mutual material attributes, exploiting commonalities and differences across varied object appearances. Such coherence enables physically sound lighting reconstruction and robust material estimation — both essential for accurate relighting. Together with a streamlined shading function and effective post-processing, ReCap outperforms all leading competitors on an expanded relighting benchmark.
Figure 1: ReCap shows a significant advantage in reconstructing environment maps with accurate tones and color fidelity. Both qualitative and quantitative assessments show that ReCap achieves more realistic and consistent relighting results under a range of unseen lighting conditions.
@@ 1 Introduction
Inspired by photometric appearance modeling [33, 25, 45] which estimates surface properties from object appearances under varied lighting, we introduce additional photometric supervision to address the albedo-lighting ambiguity. While traditional approaches rely on controlled lighting setups like light stages [11] and collocated lights [44] to provide supervision through known light directions and/or intensities, they require dedicated hardwares. Instead, we propose ReCap to leverage object captures across unknown lighting conditions, modeling light-dependent appearances with multiple environment maps that share a common Gaussian model. Conceptually, this resembles multi-task learning, where the learned environment maps act as task heads querying a shared material representation for varied object appearances. In this case, the “querying" is done by the physically-based shading function, which facilitates the separation of material and lighting. This joint optimization promotes internal consistency when accounting for varied object appearances across diverse environments, simultaneously supporting more accurate lighting reconstruction.
@@
Current relighting evaluations are limited in scope, focusing primarily on diffuse surfaces [19, 51]. For a more comprehensive assessment, we re-rendered $13$ objects from NeRF [28] and RefNeRF [40] featuring both diffuse and specular surfaces. As shown in Fig. 1, experiments on the expanded benchmark confirm the effectiveness of ReCap in producing robust and more realistic relighting results. Codes and datasets are available here.
@@
Figure 2: The proposed ReCap training framework. Compared to original 3DGS [22], each Gaussian is augmented with 3 extra material attributes. Given $k$ sets of object appearances from unknown lighting conditions as input, $k$ learnable environment maps are instantiated. Gaussian color is computed according to the shading function in the world space based on environment queries and material properties. 2D images are rasterized with standard Gaussian splatting and used for loss computation. $\mathcal{L}_{\text{image}}$: image reconstruction loss from [22]. Additional loss terms for material and geometry are not shown.
@@ 2 Related Work
Novel View Synthesis (NVS). Pure NVS techniques focus on reproducing the scene appearance under its original environment. NeRF [28] and its follow-ups [2, 3, 4] achieve remarkable NVS quality with implicit radiance fields and volume rendering. Although great progress has been made in NeRF acceleration [14, 29, 43] leveraging voxel grids or hash tables, their rendering speeds ($\sim$10 fps) are still far from interactive. Building on progress in differentiable point-based rendering [47, 23], 3DGS [22] recently introduced an explicit Gaussian representation with an efficient tile-based rasterizer, delivering photo-realistic rendering at impressive frame rates ($\sim$100 fps). This makes 3DGS well-suited for high-quality real-time relighting.
Relighting. The relighting task traditionally involves editing lighting within fixed views [12, 36, 44]. The term has since broadened to encompass novel view relighting [45, 20], which aims to generate images from both new viewpoints and under new lighting conditions. With the advent of differentiable neural rendering, reconstruction-based methods [7, 48, 50, 39] have found a promising direction to combine physically-based rendering (PBR) with neural representations for editable and realistic rendering. Closely related to our work, PBR is also incorporated by 3D Gaussians, which enable Gaussian relighting. 3DGS-DR [46] and GShader [18] both specialize in specular objects and propose customized shading functions. R3DG [15] further incorporates ray tracing for indirect illumination, while GS-IR [26] relies on baked occlusion maps. With only single environment captures as input, these Gaussian relighting methods face challenges in correctly decoupling lighting from material properties.
Albedo-lighting Disambiguation. Recovering reliable illumination and albedo from object appearances is an ill-posed problem [35] in inverse rendering. It is often simplified under the assumption of controllable lighting [6, 31], known geometry [32], or when regularized with strong priors [37, 13] in domain-specific tasks such as relighting of human faces. For more general cases, object captures from varied lighting, known as photometric images, can provide important visual cues. Several NeRF-based relighting methods have taken advantage of such captures. NeRV [39] requires multiple known lighting environments. NeRD [7] models lighting as spherical Gaussians and optimizes separate lighting for every input image. TensoIR [19] encodes lighting as an additional dimension in its factorized tensor representation and relies on ground-truth albedo scaling for accurate relighting. In contrast, we use object captures under unknown lighting, optimizing a learnable light map separately for each scene and independently from the Gaussian model, without requiring ground-truth albedo for relighting.
@@ 3.2 Disambiguate Split Sum Approximation
$\begin{split}L_{out}(\mathbf{x},\mathbf{v})&=\underbrace{E_{d}(\mathbf{n})(1-m)\mathbf{b}}_{diffuse}\\ &+\underbrace{E_{s}(\mathbf{n},\mathbf{v})\left[F_{0}\beta_{1}(r,\mathbf{n},\mathbf{v})+\beta_{2}(r,\mathbf{n},\mathbf{v})\right]}_{specular},\\ \end{split}$ (6)
@@ 3.3 Lighting Representation
While increasing the number of environments can enhance the decoupling as later shown in Sec. 4.3, it also requires additional effort to capture real objects across multiple scenes. As a proof of concept, we limit our investigation to a dual-environment setup unless mentioned otherwise.
@@
3.4 Post-shading Processing
Standard HDR maps represent radiance in linear RGB space when used as lighting sources in rendering [10]. To apply them effectively for relighting, the learned lighting values need a clear physical interpretation. Current methods adopt diverse post-shading strategies during training, influencing both the interpretation of learned environment maps and the processing of new HDR maps. For example, if the learned values are constrained to $[0,1]$ and gamma correction is applied post-shading, the environment map is effectively treated as being in a linear low dynamic range (LDR) space.
In Tab. 1, we analyze three major factors that affects relighting. Gamma correction is essential for relighting, as expected, since the shading function models linear light transport. However, it is often overlooked [15, 46], because omitting it still yields good NVS results. While previous work also consider complex tone-mappers [39] from graphics, they perform worse than simple clipping, likely due to the introduced non-linearity hindering optimization.
To balance relighting and NVS, we constrain environment maps to be positive and apply simple clipping followed by gamma correction. This implies a linear HDR lighting space where new environment maps can be used without modification, removing the need for normalizing relit images [27] or learned albedo [19] to match the ground-truth (unavailable in practice) when when evaluating relighting. Combined with cross-environment decoupling, this essentially improves relighting by (i) allowing Gaussians to push light-dependent appearances into learnable “sinks" as pseudo light maps, and (ii) regulating these “sinks" with HDR processing and PBR, enabling them to approximate real ambient light in value distribution.
@@ 3.5 Geometry Estimation
Table 2: Relighting and NVS results. Methods optimized over single-env ( ) and dual-env ( ) inputs differ in the choice of training split B. “GT scaling" refers to the use of ground truth albedo to re-scale learned albedo for relighting. Best and second best results are highlighted in bold and italic respectively.
use GT scaling Training Envs Unseen Relights Seen Relights NVS
split A split B night snow sunset interior fireplace forest Avg. bridge courtyard bridge* courtyard*
@@
3DGS-DR[46] - bridge bridge 20.18 18.99 23.19 21.33 18.76 24.90 21.22 33.18 - 35.59 -
GS-IR[26] - bridge bridge 25.56 22.44 21.22 19.46 25.64 21.10 22.57 22.01 - 31.22 -
R3DG[15] - bridge bridge 23.71 22.09 22.84 20.09 23.08 22.19 22.33 22.32 - 30.61 -
GShader[18] - bridge bridge 21.48 17.61 18.93 21.89 20.18 22.70 20.47 26.03 - 33.05 -
ReCap (ours) - bridge bridge 25.31 21.76 23.20 21.66 24.52 23.15 23.27 25.30 - 33.40 -
3DGS-DR[46] - bridge courtyard 20.53 20.25 24.56 21.71 19.41 24.25 21.78 24.89 23.92 24.84 23.57
GS-IR[26] - bridge courtyard 26.22 22.03 20.85 19.18 25.90 20.52 22.45 20.78 20.96 24.92 23.16
R3DG[15] - bridge courtyard 23.91 21.78 22.55 19.86 23.38 21.74 22.21 20.98 20.84 25.26 22.68
GShader[18] - bridge courtyard 22.73 18.20 19.96 22.11 20.97 23.02 21.17 23.34 21.75 24.96 22.84
TensoIR[19] bridge courtyard 27.22 22.15 24.31 23.12 25.35 24.80 24.49 24.50 23.82 29.13 27.46
TensoIRno scale[19] - bridge courtyard 26.24 20.11 22.75 21.79 24.45 23.29 23.11 23.50 22.47 29.13 27.46
GShader[18] + ours - bridge courtyard 23.52 17.69 19.95 23.22 21.75 24.50 21.77 24.45 24.32 31.13 29.30
ReCap (ours) - bridge courtyard 27.62 24.64 26.33 24.40 26.52 25.38 25.82 26.95 25.52 32.23 30.76
@@
3DGS-DR[46] - bridge bridge 0.882 0.894 0.925 0.907 0.843 0.906 0.893 0.971 - 0.978 -
GS-IR[26] - bridge bridge 0.902 0.904 0.890 0.861 0.884 0.867 0.885 0.898 - 0.952 -
R3DG[15] - bridge bridge 0.889 0.913 0.918 0.866 0.866 0.877 0.888 0.879 - 0.959 -
GShader[18] - bridge bridge 0.889 0.887 0.896 0.905 0.854 0.904 0.889 0.949 - 0.968 -
ReCap (ours) - bridge bridge 0.911 0.918 0.926 0.907 0.883 0.906 0.909 0.946 - 0.970 -
3DGS-DR[46] - bridge courtyard 0.885 0.902 0.934 0.907 0.846 0.903 0.896 0.930 0.927 0.926 0.924
GS-IR[26] - bridge courtyard 0.906 0.893 0.887 0.858 0.885 0.862 0.882 0.881 0.879 0.915 0.905
R3DG[15] - bridge courtyard 0.896 0.912 0.922 0.866 0.870 0.873 0.890 0.859 0.888 0.932 0.919
GShader[18] - bridge courtyard 0.904 0.891 0.908 0.906 0.862 0.905 0.896 0.925 0.913 0.928 0.922
TensoIR[19] bridge courtyard 0.908 0.861 0.891 0.870 0.886 0.888 0.884 0.893 0.883 0.962 0.957
TensoIRno scale[19] - bridge courtyard 0.910 0.861 0.893 0.871 0.889 0.889 0.885 0.895 0.884 0.962 0.957
GShader[18] + ours - bridge courtyard 0.915 0.893 0.915 0.919 0.874 0.922 0.906 0.937 0.938 0.959 0.956
ReCap (ours) - bridge courtyard 0.935 0.938 0.951 0.929 0.903 0.926 0.930 0.951 0.943 0.966 0.963
@@
3DGS-DR[46] - bridge bridge 0.081 0.091 0.073 0.085 0.108 0.083 0.087 0.041 - 0.034 -
GS-IR[26] - bridge bridge 0.099 0.094 0.101 0.119 0.108 0.109 0.105 0.099 - 0.065 -
R3DG[15] - bridge bridge 0.103 0.086 0.082 0.124 0.114 0.104 0.102 0.113 - 0.053 -
GShader[18] - bridge bridge 0.091 0.116 0.098 0.094 0.113 0.091 0.100 0.062 - 0.044 -
ReCap (ours) - bridge bridge 0.077 0.080 0.073 0.089 0.092 0.084 0.083 0.061 - 0.042 -
3DGS-DR[46] - bridge courtyard 0.084 0.096 0.078 0.096 0.115 0.094 0.094 0.078 0.080 0.082 0.085
GS-IR[26] - bridge courtyard 0.101 0.107 0.107 0.127 0.112 0.116 0.112 0.115 0.119 0.099 0.104
R3DG[15] - bridge courtyard 0.101 0.092 0.083 0.129 0.117 0.110 0.105 0.127 0.106 0.078 0.089
GShader[18] - bridge courtyard 0.087 0.115 0.095 0.101 0.114 0.097 0.101 0.086 0.094 0.080 0.086
TensoIR[19] bridge courtyard 0.131 0.155 0.134 0.138 0.138 0.127 0.137 0.128 0.133 0.093 0.098
TensoIRno scale[19] - bridge courtyard 0.133 0.159 0.136 0.141 0.138 0.131 0.140 0.130 0.137 0.093 0.098
GShader[18] + ours - bridge courtyard 0.074 0.113 0.084 0.089 0.102 0.080 0.090 0.073 0.073 0.055 0.059
ReCap (ours) - bridge courtyard 0.059 0.069 0.057 0.077 0.075 0.071 0.068 0.058 0.064 0.047 0.051
@@ 4.1 Implementation Details
Datasets. We construct a more comprehensive RelightObj dataset by relighting 8 general objects from NeRF Synthetic Dataset [28] and 5 shiny objects from the Shiny Blender Dataset [40] under 8 different HDR maps. Each scene includes 200 training views and 200 test views, with identical camera poses across scenes. To eliminate biases from training view poses, the training views are divided into two splits, A and B, each containing 100 views. In the single-environment (single-env) setup, both splits come from the same environment, whereas in the dual-environment (dual-env) setup, split B is taken from a different environment. Relighting performance is judged on unseen environments.
@@
Baseline and Metrics. We compare with the following baselines: (a) GShader [18]: a Gaussian rendering method utilizing a customized shading function with a residual color term; (b) GS-IR [26]: an inverse rendering approach using view-space shading with a microfacet BRDF model and a baked occlusion map; (c) 3DGS-DR [46]: a Gasussian rendering method targeting reflective surfaces with a customized shading function and normal propagation; (d) R3DG [15]: a Gaussian relighting technique that incorporates ray tracing and a microfacet BRDF model; and (e) TensoIR [19]: top-performing NeRF-based method in novel view synthesis and relighting. Following these works, rendering and relighting quality are evaluated using PSNR, SSIM [42], and LPIPS [49].
@@
Training and Testing. All experiments are conducted on an Nvidia RTX 3090 graphics card. Our models are optimized using Adam for 30,000 iterations. All other methods are retrained on our dataset using their official repositories and recommended settings, with results reported from the retrained models. As GShader [18] and 3DGS-DR [46] lack official relighting code, we pre-processed HDR maps following their design choices (details in the supplementary). For TensoIR [19], which uses per-object hyperparameters, we tested all provided settings and report the best results.
@@
Figure 6: Qualitative comparison of NVS and relighting results. Upper: toaster, lower: chair. TensoIR produces accurate relighting with GT albedo scaling for chair, but suffers from tone-shift without it. For our method, specular surfaces in toaster provide more information for lighting reconstruction, leading to better albedo-lighting decoupling hence more truthful relighting results than mostly diffuse chair. Refer to the supplementary for more visuals.